OpenStudy (anonymous):

multiply (5-3i)(6+4i)

OpenStudy (anonymous):

You need to Foil : First, Inside, Outside, Last First- 5*6=30 Inside- (-3i)*6=-18i Outside- 4i*5=20i Last- (-3i)*(4i)=(-12)(-1) because i^2 is -1 Then add them all together 20-18i+20i+12=2i+32

OpenStudy (anonymous):

(5-3i)(6+4i)=5(6+4i)-3i(6+4i)=30+20i-18i-12(i^2)=30+2i-12(-1)=42+2i

OpenStudy (anonymous):

oh yeah I wrote 20 instead of 30 in the last line.